Select cost information for Seacrest Enterprises is as follows:

1,000 units of output

5,000 units of output

Total

Cost/Unit

Total

Cost/Unit

Direct materials

$

5,000

$

5.00

$

25,000

$

5.00

Utilities expense

$

1,000

$

1.00

$

3,750

$

0.75

Rent expense

$

4,000

$

4.00

$

4,000

$

0.80

Based on this information it has been determined that Utilities Expense is a mixed cost and rent expense is a fixed cost while direct materials is variable cost. Explain in narrative form why this is true and be sure to provide definitions of what exactly mixed, fixed and variable costs are. Also, include in your answer a definition and illustration of step-wise cost.
Answer #1

Variable cost are the cost based on the number of units to be produced. They change based on the level of production. Variable cost per unit remains the same.

Fixed cost are the cost that do not change with change in level of activity of the company. They are fixed and thus per unit fixed cost reduces when more production is done.

Mixed cost are the cost that have element of both the cost i.e Variable & Fixed.

1000 Units 5000 Units
Total Cost/Unit Total Cost/Unit
Direct materials 5000 5 25000 5
Utilities expense 1000 1 3750 0.75
Rent expense 4000 4 4000 0.8

It can be seen from the above table that per unit Direct Material cost is same in both the level of activities. Thus it is Variable Cost

Rent Expense is unchanged irrespective of the change in number of units produced. Thus this is Fixed Cost

Rent Expense is Mixed Cost. Variable & Fixed Component can be identified as follows:

Change in Utilities Expense = 3750 - 1000 = 2750

Change in Output (Units) = 5000 - 1000 = 4000

Variable Cost per Unit = 2750 / 4000 = 0.6875 per unit

Fixed Portion of Utilities = 3750 - (5000*0.6875) = 312.50